# On the representation of stack operators by mathematical morphology

### v1

@@ Abstract
This paper introduces the class of grey-scale image stack operators as those that (a) map binary-images into binary-images and (b) commute in average with cross-sectioning. We show that stack operators are 1-Lipchitz extensions of set operators which can be represented by applying a characteristic set operator to the cross-sections of the image and summing. In particular, they are a generalisation of stack filters, for which the characteristic set operators are increasing. Our main result is that stack operators inherit lattice properties of the characteristic set operators. We focus on the case of translation-invariant and locally defined stack operators and show the main result by deducing the characteristic function, kernel, and basis representation of stack operators. The results of this paper have implications on the design of image operators, since imply that to solve some grey-scale image processing problems it is enough to design an operator for performing the desired transformation on binary images, and then considering its extension given by a stack operator. We leave many topics for future research regarding the machine learning of stack operators and the characterisation of the image processing problems that can be solved by them.
@@ 1 Introduction
Stack filters, proposed by [13] in the 80’s, is a classical family of filters for signal and image processing. Frameworks for learning stack filters from data have been extensively studied [3, 4, 6] and many algorithms to design stack filters have been proposed [5, 14]. In the context of mathematical morphology, properties of stack filters have been deduced in two important works by Maragos and Schafer [7, 8].
@@
for all $f\in\{0,\dots,m\}^{E}$, in which $T_{t}$ is the cross-section of $f$ at point $t$. We refer to [7, 8] for a proof of this fact. In this paper, we study the class of stack operators that are the operators $\psi$ that satisfy (1) for a not necessarily increasing set operator $\tilde{\psi}$.
@@
We define the stack operators as those that (a) map binary functions into binary functions and (b) such that $\psi(f)$ equals the mean of $\psi(mT_{t}[f])$ over $t=1,\dots,m$, and then show that these operators can be written as (1). In particular, we show that stack operators are an extension of set operators (i.e, binary image operators) to function operators (i.e., grey-scale image operators). Our main result is that the stack operator $\psi$ inherits many properties of the respective set operator $\tilde{\psi}$, such as being an erosion, a dilation, an anti-erosion, an anti-dilation, increasing, decreasing, extensive, anti-extensive, sup-generating and inf-generating. For simplicity, we focus on the case of translation-invariant and locally defined stack operators and show the main result by deducing the characteristic function, kernel and basis representation of stack operators.
@@
The results of this paper have implications on the design of image operators, since allow to solve grey-scale image processing tasks by designing an operator for performing the desired transformation on binary images and then considering its extension given by a stack operator. We also discuss how the results can be leveraged to develop learning methods for grey-scale image processing based on deep neural networks and discrete morphological neural networks [9].
@@
In Section 2, we define the concepts necessary to introduce the stack operators in Section 3, where we show that the space of stack operators is isomorphic to the space of set operators, and that stack operators are 1-Lipschitz extensions of set operators. In Section 4, we show that the stack filters are the increasing stack operators. In Section 5, we establish the characteristic function, kernel and basis representation of stack W-operators. In Section 6, we state and prove the main result of this paper establishing that stack W-operators inherit lattice properties of the associated set operator, and we discuss the implications of our results to the design of image operators. In Section 7, we give our final remarks. Proofs for the main results are presented in the Appendix 0.A.
2 Grey-scale images and set operators
@@
Let $E$ be a countable set, $(E,+)$ be an Abelian group, with zero element $o\in E$, and $M=\{0,\dots,m\}$ for a positive integer $m$ fixed. A function $f:E\mapsto M$ is a grey-scale image, in which $f(x)$ represents the intensity of the pixel $x\in E$. In particular, a grey-scale binary image is such that $f(x)\in\{0,m\}$ for all $x\in E$. The set $M^{E}$ of all such images is equipped with a partial order satisfying $f,g\in M^{E},f\leq g\iff f(x)\leq g(x),\forall x\in E$. The image threshold at level $t$ is a mapping $T_{t}:M^{E}\mapsto\{0,1\}^{E}$ given by, for any $f\in M^{E},t\in M$ and $x\in E$, $T_{t}[f](x)=\mathds{1}\{f(x)\geq t\}$. The binary function $T_{t}[f]$ is called the cross-section of $f$ at level $t$. The collection $\{T_{t}:t\in M\}$ has some simple, yet important, properties, that we state without proof as a proposition.
Proposition 1
The collection of image threshold mappings $\{T_{t}:t\in M\}$ satisfies
  1. (i) For all $f\in M^{E}$ and $x\in E$, $f(x)=\sum_{t=1}^{m}T_{t}[f](x)$.
  2. (ii) Let $t_{1},t_{2}\in M$ and $f\in M^{E}$. If $t_{1}\leq t_{2}$, then $T_{t_{2}}[f]\leq T_{t_{1}}[f]$.
  3. (iii) For $f,g\in M^{E}$, $f\leq g\iff T_{t}[f]\leq T_{t}[g]$ for all $t\in M$.
An image operator is a mapping $\psi:M^{E}\mapsto M^{E}$ and the set of all such operators, denoted by $\Psi_{M}$, is equipped with a partial order satisfying $\psi_{1},\psi_{2}\in\Psi_{M},\psi_{1}\leq\psi_{2}\iff\psi_{1}(f)\leq\psi_{2}(f),\forall f\in M^{E}$. The poset $(\Psi_{M},\leq)$ is a complete lattice.
@@
A binary image can be represented as a function $X:E\mapsto\{0,1\}$ in which $X(x)=1$ if, and only if, $x\in E$ is a pixel in the image foreground. The set $\{0,1\}^{E}$ of all binary images equipped with the partial order satisfying $X,Y\in\{0,1\}^{E},X\leq Y\iff X(x)\leq Y(x),\forall x\in E$, is a complete lattice. Mappings from $\{0,1\}^{E}$ into itself are called set operators, a nomenclature inspired by the fact that $\{0,1\}^{E}$ is isomorphic to the power-set of $E$. We denote the space of all set operators by $\Psi$ and equip it with the point-wise partial order $\tilde{\psi}_{1},\tilde{\psi}_{2}\in\Psi,\tilde{\psi}_{1}\leq\tilde{\psi}_{2}\iff\tilde{\psi}_{1}(X)\leq\tilde{\psi}_{2}(X),\forall X\in\{0,1\}^{E}$. An image operator is an extension of a set operator if they coincide on binary images.
@@
For each $h\in E$, denote by $\tau_{h}:E\mapsto E$ the translation by $h$ function that maps $x\mapsto x+h$. An operator $\psi\in\Psi_{M}$, and a set operator $\tilde{\psi}\in\Psi$, are translation-invariant if they commute with translation, that is, $\psi(f\circ\tau_{h})=\psi(f)\circ\tau_{h}$ and $\tilde{\psi}(X\circ\tau_{h})=\tilde{\psi}(X)\circ\tau_{h}$ for all $f\in M^{E}$, $X\in\{0,1\}^{E}$ and $h\in E$.
For $W\in\{0,1\}^{E}$, denote by $X|_{W}$ and $f|_{W}$ the restriction of $X\in\{0,1\}^{E}$ and $f\in M^{E}$ to the set $\chi(W)\coloneqq\{x\in E:W(x)=1\}$. An operator $\psi\in\Psi_{M}$, and a set operator $\tilde{\psi}\in\Psi$, are locally defined within $W$ if $\psi(f)(x)=\psi(f|_{W\circ\tau_{-x}})(x)$ and $\tilde{\psi}(X)(x)=\tilde{\psi}(X|_{W\circ\tau_{-x}})(x)$ for all $f\in M^{E}$, $X\in\{0,1\}^{E}$ and $x\in E$. The locally defined condition implies, for example, that the value of the transformed image $\psi(f)$ at point $x$ depends only on the value of $f$ in the neighbourhood $\chi(W\circ\tau_{-x})=\{h\in E:W(h-x)=1\}$ of $x$. Operators that are translation-invariant and locally defined are called W-operators. The collections of image and set W-operators are denoted by $\Psi_{M}^{W}\subseteq\Psi_{M}$ and $\Psi^{W}\subseteq\Psi$, respectively.
@@ 3 Stack operators
An image operator $\psi\in\Psi_{M}$ is a stack operator if, and only if,
  • (a) For all $x\in E$ and $X\in\{0,1\}^{E}$, $\psi(mX)(x)\in\{0,m\}$.
  • (b) For all $f\in M^{E}$
$\psi(f)=\frac{1}{m}\sum_{t=1}^{m}\psi(mT_{t}[f]).$ (2)
Condition (a) of Definition 2 implies that stack operators take grey-scale binary images $mX\in\{0,m\}^{E}$ into grey-scale binary images. It follows from this condition that $m^{-1}\psi(mT_{t}[f])\in\{0,1\}^{E}$ and therefore the right-hand side of (2) is an element of $M^{E}$. Condition (b) then implies that $\psi(f)$ is the average of the transformation by $\psi$ of the cross-sections of $f$ represented as grey-scale binary images. Stack operators have an associated set operator $\tilde{\psi}:\{0,1\}^{E}\mapsto\{0,1\}^{E}$ and the collection $\Psi$ of set operators is lattice isomorphic to the collection of stack image operators $\Sigma_{S}$. In particular, $(\Sigma_{S},\leq)$ is a complete lattice. We state without proof this result that is a direct consequence of the definitions of the isomorphisms $\mathcal{F}$ and $\mathcal{F}^{-1}$.
@@
We call $\mathcal{F}^{-1}(\psi)$ the characteristic set operator of $\psi\in\Sigma_{S}$. It follows from Proposition 2 that $m^{-1}\psi(mX)=\mathcal{F}^{-1}(\psi)(X)$ and therefore $\psi$ is an extension of its characteristic set operator (ref. Definition 1). Moreover, this extension is 1-Lipschitz considering the $L^{1}$-norm in the domain of $\psi$ and the $L^{\infty}$-norm in its image, that is:
$\lVert\psi(f)-\psi(g)\rVert_{\infty}\leq\lVert f-g\rVert_{1},$ $\psi\in\Sigma_{S},f,g\in M^{E}.$ (3)
In particular, stack operators are $(L^{1},L^{\infty})$-continuous. A proof for the next corollary is in Appendix 0.A.
@@
Remark 1
The inequality (3) is tight since, for example, equality is attained when $\psi$ is the identity operator and $g=f+s\mathds{1}_{\{x\}}$ for $s\in\{1,\dots,m\}$ and $x\in E$ fixed.
An important class of stack operators is that of the stack W-operators, denoted by $\Sigma_{S}^{W}=\Psi_{M}^{W}\cap\Sigma_{S}$. The isomorphism $\mathcal{F}$ preserves the W-operator properties, hence maps set W-operators to stack W-operators. We state without proof this result that is a direct consequence of the definitions of $\mathcal{F}$ and $\mathcal{F}^{-1}$.
@@
In Figure 1 we present two simple examples of stack operators: for boundary recognition and noise filtering. In these cases we can clearly see that the stack operators extend to grey-scale images the transformation performed on set operators. In the example of boundary recognition, the operator is not increasing, therefore is not a stack filter.
Figure 1: (a) Binary and grey-scale input images. (b) Binary and grey-scale input images with 2.5% salt and pepper noise. (c) Boundary recognition of (a) by the set, and associated stack, operator with 3 x 3 window and characteristic function $\tilde{\phi}$ with $\tilde{\phi}(x)=0$ if, and only if, $x=\boldsymbol{0}$ or $x=\boldsymbol{1}$. (d) Noise filtering of (b) by an alternating sequential set, and associated stack, filter with 3 x 3 structuring element. (e) Result obtained by composing to (b) the operators in (c) and (d). All images are 256 x 256 pixels and $m=255$.
4 Stack filters
The stack filters are operators $\psi$ that commute with threshold in the following sense:
@@ 4 Stack filters
for all $t\geq 1$ and $f\in M^{E}$. Equivalently, it follows from the definition of $\mathcal{F}^{-1}$ that a stack operator $\psi$ with characteristic set operator $\tilde{\psi}$ is a stack filter if, and only if, $T_{t}[\psi(f)]=\tilde{\psi}(T_{t}[f])$ for all $t\geq 1$ and $f\in M^{E}$. A stack filter is a stack operator since it follows from (4) that $\psi(mX)(x)=mT_{1}[\psi(X)](x)\in\{0,m\}$ for all $x\in E$ and $X\in\{0,1\}^{E}$ and that $\psi(f)=\sum_{t=1}^{m}T_{t}[\psi(f)]=\frac{1}{m}\sum_{t=1}^{m}\psi(mT_{t}[f])$ in which the first equality follows from Proposition 1 (i) and the second from (4). The singularity of stack filters is that they satisfy (4) for all $t\geq 1$ while general stack operators satisfy (4) summing over $t$ (cf. (2)). Actually, the stack filters are the stack operators with increasing characteristic set operator. It follows from Proposition 4 that the definition of stack filters by equation (4) is equivalent to the classical definition of stack filters [7]. The proof of this proposition is analogous to that of [7, Theorem 3].
@@
An operator $\psi\in\Sigma_{S}$ is a stack filter if, and only if, $\mathcal{F}^{-1}(\psi)$ is increasing.

### v2

@@ Abstract
This paper introduces the class of grey-scale image stack operators as those that (a) map binary-images into binary-images and (b) commute on average with cross-sectioning. Equivalently, stack operators are 1-Lipchitz extensions of set operators which can be represented by applying a characteristic set operator to the cross-sections of the image and adding. In particular, they are a generalisation of stack filters, for which the characteristic set operators are increasing. Our main result is that stack operators inherit lattice properties of the characteristic set operators. We focus on the case of translation-invariant and locally defined stack operators and show the main result by deducing the characteristic function, kernel, and basis representation of stack operators. The results of this paper have implications on the design of image operators, since imply that to solve some grey-scale image processing problems it is enough to design an operator for performing the desired transformation on binary images, and then considering its extension given by a stack operator. We leave many topics for future research regarding the machine learning of stack operators and the characterisation of the image processing problems that can be solved by them.
@@ 1 Introduction
Stack filters, proposed by [17] in the 80s, is a classical family of filters for signal and image processing. Frameworks for learning stack filters from data have been extensively studied [4, 5, 7] and many algorithms to design stack filters have been proposed [6, 18]. In the context of mathematical morphology, properties of stack filters have been deduced in two important works by Maragos and Schafer [9, 10].
@@
for all $f\in\{0,\dots,m\}^{E}$, in which $T_{t}[f](x)=\mathds{1}\{f(x)\geq t\},x\in E,$ is the cross-section of $f$ at level $t\in\{0,\dots,m\}$. We refer to [9, 10] for a proof of this fact.
@@
In this paper, we study the class of stack operators that are the operators $\psi$ that satisfy (1) for a not necessarily increasing set operator $\tilde{\psi}$. We show that stack operators can be equivalently defined as those that map binary functions into binary functions and satisfy
$\psi(f)=\frac{1}{m}\sum_{t=1}^{m}\psi(mT_{t}[f]),$ (2)
that is, $\psi(f)$ is the average of $\psi$ applied to the cross-sections of $f$ rescaled to a function taking values in $\{0,m\}$. We interpret (2) as $\psi$ commuting on average with cross-sectioning. In particular, we show that stack operators are an extension of set operators (i.e, binary image operators) to function operators (i.e., grey-scale image operators).
@@
Our main result is that the stack operator $\psi$ inherits many properties of the respective set operator $\tilde{\psi}$, such as being an erosion, a dilation, an anti-erosion, an anti-dilation, increasing, decreasing, extensive, anti-extensive, sup-generating and inf-generating. For simplicity, we focus on the case of translation-invariant and locally defined stack operators and show the main result by deducing the characteristic function, kernel, and basis representation of stack operators.
@@
The results of this paper have implications on the design of image operators, since allow to solve grey-scale image processing tasks by designing an operator for performing the desired transformation on binary images and then considering its extension given by a stack operator via (1). In particular, the results can be leveraged to develop learning methods for grey-scale image processing based on deep neural networks and discrete morphological neural networks [11].
We consider two examples of stack operators designed via (1), one for image denoising and another for boundary recognition, both acting on functions $f:\mathbb{Z}^{2}\mapsto\{0,\dots,m\}$ and with associated set operators acting on functions $X:\mathbb{Z}^{2}\mapsto\{0,1\}$. The denoising stack operator, denoted by $\psi_{1}$, is the extension of the alternating-sequential filter $\tilde{\psi}_{1}=\epsilon_{W}\,\delta_{W}\,\delta_{W}\,\epsilon_{W}$, in which $W$ is the 3 x 3 square centred at the origin of $\mathbb{Z}^{2}$, and $\epsilon_{W}$ and $\delta_{W}$ are the set erosion and dilation with structuring element $W$, respectively. This set operator is obtained by applying the opening $\delta_{W}\,\epsilon_{W}$ followed by the closing $\epsilon_{W}\,\delta_{W}$, and is known to perform denoising. Since this set operator is increasing, the associated stack operator is actually a stack filter. An example of the action of this set operator, and the associated stack operator, is presented in Figure 1 (d). Observe that the stack filter $\psi_{1}$ inherits the denoising property of $\tilde{\psi}_{1}$.
The boundary recognition stack operator, denoted by $\psi_{2}$, is obtained from a translation-invariant and locally defined set operator [^1] See Section 2 for a formal definition of translation-invariant and locally defined set operators. satisfying $\tilde{\psi}_{2}(X)(x)=0$ if $X(y_{1})=X(y_{2})\ \forall y_{1},y_{2}\in x+W$ and equals $1$ otherwise, for $x\in\mathbb{Z}^{2}$, in which $x+W$ is the 3 x 3 square centred at $x$. In other words, $\tilde{\psi}_{2}(X)(x)$ equals 0 if all pixels of $X$ in a neighbourhood of $x$ have the same value, and equals $1$ otherwise. We note that this set operator is not increasing, so the associated stack operator is not a stack filter. Furthermore, as can be seen in the example in Figure 1 (c), this set operator has the effect of boundary recognition, a property that is inherited by the associated stack operator. In Figure 1 (e) we present the result obtained by applying $\tilde{\psi}_{2}\tilde{\psi}_{1}$ and $\psi_{2}\psi_{1}$, that is, boundary recognition after denoising. Even though we focus on these examples, the class of stack operators encompasses other operators of interest, such as grey-scale top hat transforms, and morphological gradients, with flat structuring elements.
Figure 1: (a) Binary and grey-scale input images. (b) Binary and grey-scale input images with 2.5% salt and pepper noise. (c) Boundary recognition of (a) by the set operator $\tilde{\psi}_{2}$ defined in Section 1, and associated stack operator $\psi_{2}$. (d) Noise filtering of (b) by the alternating-sequential filter $\tilde{\psi}_{1}=\epsilon_{W}\,\delta_{W}\,\delta_{W}\,\epsilon_{W}$, and the associated stack filter $\psi_{1}$. (e) Result obtained by applying $\tilde{\psi}_{2}\tilde{\psi}_{1}$ and $\psi_{2}\psi_{1}$ to the images in (b). All images are 256 x 256 pixels and $m=255$.
@@
In Section 2, we define the concepts necessary to introduce the stack operators in Section 3. In Section 4, we establish the characteristic function, kernel, and basis representation of stack W-operators. In Section 5, we state and prove the main result of this paper, and we discuss the implications of our results to the design of image operators. In Section 6, we give directions for future research. Proofs for the main results are presented in the Appendix 0.A.
2 Grey-scale image and set operators
@@
Let $E$ be a countable set, $(E,+)$ be an Abelian group, with zero element $o\in E$, and $M=\{0,\dots,m\}$ for a positive integer $m$ fixed. A function $f:E\mapsto M$ is a grey-scale image, in which $f(x)$ represents the intensity of the pixel $x\in E$. In particular, a grey-scale binary image is such that $f(x)\in\{0,m\}$ for all $x\in E$. The set $M^{E}$ of all grey-scale images is equipped with a partial order satisfying $f,g\in M^{E},f\leq g\iff f(x)\leq g(x),\forall x\in E$. The image cross-section, or threshold, at level $t$ is a mapping $T_{t}:M^{E}\mapsto\{0,1\}^{E}$ given by, for any $f\in M^{E},t\in M$ and $x\in E$, $T_{t}[f](x)=\mathds{1}\{f(x)\geq t\}$. An image operator is a mapping $\psi:M^{E}\mapsto M^{E}$ and the set of all such operators, denoted by $\Psi_{M}$, is equipped with a partial order satisfying $\psi_{1},\psi_{2}\in\Psi_{M},\psi_{1}\leq\psi_{2}\iff\psi_{1}(f)\leq\psi_{2}(f),\forall f\in M^{E}$. The poset $(\Psi_{M},\leq)$ is a complete lattice.
@@
A binary image can be represented as a function $X:E\mapsto\{0,1\}$ in which $X(x)=1$ if, and only if, $x\in E$ is a pixel in the image foreground. The set $\{0,1\}^{E}$ of all binary images equipped with the partial order satisfying $X,Y\in\{0,1\}^{E},X\leq Y\iff X(x)\leq Y(x),\forall x\in E$, is a complete lattice. Mappings from $\{0,1\}^{E}$ into itself are called set operators, a nomenclature inspired by the fact that $\{0,1\}^{E}$ is isomorphic to the power-set of $E$. We denote the space of all set operators by $\Psi$ and equip it with the point-wise partial order $\tilde{\psi}_{1},\tilde{\psi}_{2}\in\Psi,\tilde{\psi}_{1}\leq\tilde{\psi}_{2}\iff\tilde{\psi}_{1}(X)\leq\tilde{\psi}_{2}(X),\forall X\in\{0,1\}^{E}$. An image operator is an extension of a set operator if they coincide on binary images. We note that the denoising and boundary recognition operators $\psi_{1}$ and $\psi_{2}$ introduced in Section 1 are extensions of the respective set operator.
@@
For each $h\in E$, denote by $\tau_{h}:E\mapsto E$ the translation by $h$ function that maps $x\mapsto x+h$. An operator $\psi\in\Psi_{M}$, and a set operator $\tilde{\psi}\in\Psi$, are translation-invariant if they commute with translation, that is, $\psi(f\circ\tau_{h})=\psi(f)\circ\tau_{h}$ and $\tilde{\psi}(X\circ\tau_{h})=\tilde{\psi}(X)\circ\tau_{h}$ for all $f\in M^{E}$, $X\in\{0,1\}^{E}$ and $h\in E$, in which $\circ$ means function composition.
For $W\in\{0,1\}^{E}$, denote by $X|_{W}$ and $f|_{W}$ the restriction of $X\in\{0,1\}^{E}$ and $f\in M^{E}$ to the set $\chi(W)\coloneqq\{x\in E:W(x)=1\}$. An operator $\psi\in\Psi_{M}$, and a set operator $\tilde{\psi}\in\Psi$, are locally defined within $W$ if [^2] The restriction $f|_{W\circ\tau_{-x}}$ should be understood as being equal to $f(y)$ for $y\in\chi(W\circ\tau_{-x})$ and equal to 0 for $y$ not in this set, so it is an element of $M^{E}$ and $\psi(f|_{W\circ\tau_{-x}})$ is well defined. By an analogous reason, $\tilde{\psi}(X|_{W\circ\tau_{-x}})$ is well defined. $\psi(f)(x)=\psi(f|_{W\circ\tau_{-x}})(x)$ and $\tilde{\psi}(X)(x)=\tilde{\psi}(X|_{W\circ\tau_{-x}})(x)$ for all $f\in M^{E}$, $X\in\{0,1\}^{E}$ and $x\in E$. The locally defined condition implies, for example, that the value of the transformed image $\psi(f)$ at point $x$ depends only on the value of $f$ in the neighbourhood $\chi(W\circ\tau_{-x})=\{h\in E:W(h-x)=1\}$ of $x$. Operators that are translation-invariant and locally defined are called W-operators. The collections of image and set W-operators are denoted by $\Psi_{M}^{W}\subseteq\Psi_{M}$ and $\Psi^{W}\subseteq\Psi$, respectively. We note that the example operators $\psi_{1}$ and $\psi_{2}$ are W-operators, the latter by construction and the former since erosions and dilations are W-operators, and composition preserves the W-operator properties, but in a greater window [^3] See [3, Corollary 5.1] for a proof of this fact for set operators. That it also holds for stack operators follows from Proposition 2..
@@ 3 Stack operators
An image operator $\psi\in\Psi_{M}$ is a stack operator if, and only if, there exists a set operator $\tilde{\psi}\in\Psi$ such that (1) holds for all $f\in M^{E}$, or equivalently, if, and only if, (a) for all $x\in E$ and $X\in\{0,1\}^{E}$, $\psi(mX)(x)\in\{0,m\}$ and (b) relation (2) holds for all $f\in M^{E}$.
Condition (a) of Definition 2 implies that stack operators take grey-scale binary images $mX\in\{0,m\}^{E}$ into grey-scale binary images. It follows from this condition that $m^{-1}\psi(mT_{t}[f])\in\{0,1\}^{E}$ and therefore the right-hand side of (2) is an element of $M^{E}$. Condition (b) then implies that $\psi(f)$ is the average of the transformation by $\psi$ of the cross-sections of $f$ represented as grey-scale binary images. Equivalently, stack operators have an associated set operator $\tilde{\psi}:\{0,1\}^{E}\mapsto\{0,1\}^{E}$ satisfying (1), and the collection $\Psi$ of set operators is lattice isomorphic to the collection of stack image operators $\Sigma_{S}$. In particular, $(\Sigma_{S},\leq)$ is a complete lattice. We state without proof this result that is a direct consequence of the definitions of $\mathcal{F}$ and $\mathcal{F}^{-1}$ below. This proposition implies that the definitions of stack operators in Definition 2 are indeed equivalent.
@@
We call $\mathcal{F}^{-1}(\psi)$ the characteristic set operator of $\psi\in\Sigma_{S}$. It follows from Proposition 1 that $m^{-1}\psi(mX)=\mathcal{F}^{-1}(\psi)(X)$ and therefore $\psi$ is an extension of its characteristic set operator. Moreover, this extension is 1-Lipschitz considering the $L^{1}$-norm in the domain of $\psi$ and the $L^{\infty}$-norm in its image, that is [^4] Observe that the Lipschitz inequality is tight since, for example, equality is attained when $\psi$ is the identity operator and $g=f+s\mathds{1}_{\{x\}}$ for $s\in\{1,\dots,m\}$ and $x\in E$ fixed.: $\lVert\psi(f)-\psi(g)\rVert_{\infty}\leq\lVert f-g\rVert_{1}$ for $\psi\in\Sigma_{S},f,g\in M^{E}$. In particular, stack operators are $(L^{1},L^{\infty})$-continuous. A proof for the next corollary is in Appendix 0.A.
@@
An important class of stack operators is that of the stack W-operators, denoted by $\Sigma_{S}^{W}=\Psi_{M}^{W}\cap\Sigma_{S}$. The isomorphism $\mathcal{F}$ preserves the W-operator properties, and hence maps set W-operators to stack W-operators. We state without proof this result that is a direct consequence of the definitions of $\mathcal{F}$ and $\mathcal{F}^{-1}$. Clearly, the example operators $\psi_{1}$ and $\psi_{2}$ inherit the W-operator properties of their characteristic set operators.
@@
The stack filters have the property of commuting with threshold in the following sense:
@@
for all $t\geq 1$ and $f\in M^{E}$. Indeed, it follows from the definition of $\mathcal{F}^{-1}$ that a stack operator $\psi$ with characteristic set operator $\tilde{\psi}$ satisfies (3) if, and only if, $T_{t}[\psi(f)]=\tilde{\psi}(T_{t}[f])$ for all $t\geq 1$ and $f\in M^{E}$, a condition that is known to equivalently define stack operators (see [7, Definition 2.1]).
The singularity of stack filters is that they satisfy (3) for all $t\geq 1$ while general stack operators satisfy (3) summing over $t$ (see (2)), and this fact actually implies that the stack filters are the stack operators with increasing characteristic set operator. We state this fact as a proposition whose proof is analogous to [9, Theorem 3].
@@
A stack operator $\psi\in\Sigma_{S}$ is a stack filter if, and only if, $\mathcal{F}^{-1}(\psi)$ is increasing.